OpenStudy (31356):

Factoring Using the Distributive Property 2y+6xy Show steps to find GCF MEDAL REWARDED!!

OpenStudy (31356):

@jim_thompson5910

OpenStudy (mertsj):

Factor out 2y

OpenStudy (31356):

1*2y

jimthompson5910 (jim_thompson5910):

6xy = 2y*_____

OpenStudy (31356):

3xy or 3x?

jimthompson5910 (jim_thompson5910):

well if it was 3xy, then 6xy = 2y*3xy 6xy = (2*3)*x*(y*y) 6xy = 6xy^2 ... which is NOT an identity

jimthompson5910 (jim_thompson5910):

so it has to be 3x

OpenStudy (31356):

Okay thanks!

jimthompson5910 (jim_thompson5910):

so 2y+6xy becomes 2y*1 + 2y*3x I'm sure you know what to do from here

OpenStudy (mertsj):